# Be More Diverse than the Most Diverse: Online Selection of Diverse Mixtures of Generative Models

### v1

@@ 6 Numerical Results
  • Mixture Oracle. In the mixture oracle algorithm (Section 5), an oracle tells us the optimal mixture $\boldsymbol{\alpha}^{*}$ in advance, and we pull arms randomly according to this distribution. The optimal mixture is calculated by solving the quadratic optimization in Section 4 on a large number of samples for each arm. The number of chosen samples varies based on the experiments. This is an unrealistic setting that only serves as a theoretical upper bound of the performance of any online algorithm. A realistic algorithm that performs close to the mixture oracle would be almost optimal.
@@
  • Successive Halving. The Success Halving algorithm [1, 2, 3] which serves as a second baseline for comparison.
@@
We use DINOv2-ViT-L/14 [33] for image feature extraction, as recommended by [34], and use RoBERTa [35] as text encoder. Detailed explanation of the setup for each experiment is presented in Section C.1.
@@ 6.1 Optimal Mixture for Diversity and Quality via KID
We conduct three experiments to evaluate our method using the Kernel Inception Distance (KID) metric. In the first experiment, we use five distinct generative models: LDM [36], StyleGAN-XL [37], Efficient-vdVAE [38], InsGen [39], and StyleNAT [40], all trained on the FFHQ dataset [41]. In the second experiment, we use generated images from four models [^7] FFHQ and LSUN-Bedroom datasets were downloaded from the dgm-eval repository [34] (licensed under MIT license): https://github.com/layer6ai-labs/dgm-eval.: StyleGAN [41], Projected GAN [42], iDDPM [43], and Unleashing Transformers [44], all trained on the LSUN-Bedroom dataset [45]. This experiment follows a similar setup to the first. In the final experiment, we employ the Truncation Method [46, 41] to generate diversity-controlled images centered on eight randomly selected points, using StyleGAN2-ADA [47], also trained on the FFHQ dataset. Figure 3 demonstrates that the mixture of generators achieves a better KID score compared to individual models. Additionally, the two Mixture-UCB algorithms consistently outperform the baselines.
@@ 6.2 Optimal Mixture for Diversity via RKE
We use the RKE Mode Count [7] as an evaluation metric to show the effect of mixing the models on the diversity and the advantage of our algorithms Mixture-UCB-CAB and Mixture-UCB-OGD. The score in the plots is the RKE Mode Count, written as RKE for brevity.
(a) FFHQ truncated generators
(b) AFHQ truncated generators
Figure 4: Performance comparison of online algorithms based on the RKE metric for Simulator Unconditional Generative Models.
@@
We conduct two experiments on diversity-limited generative models. First, we use eight center points with a truncation value of 0.3 to generate images using StyleGAN2-ADA, trained on the FFHQ dataset. In the second experiment, we apply the same model, trained on the AFHQ Cat dataset [48], with a truncation value of 0.4. As shown in Figure 4, the mixture achieves a higher RKE score, and our algorithms consistently get a higher RKE value. The increase in diversity is visually depicted in Figures 8 and 9.
@@
Figure 5: Visual comparison of the diversity across individual arms and the optimal mixture for Dog Breed Generators and Style-Specific Generators.
Given that many text-to-image generative models are trained on different datasets, we expect a more pronounced increase in diversity when using a mixture of these models, compared to unconditional generative models trained on the same dataset, such as FFHQ.
We explore three scenarios to evaluate the performance of our algorithms. First, we simulate when models struggle to capture different styles. To address this, we use Stable Diffusion XL [49] with specific prompts to create three car image generators with distinct styles: realistic, surreal, and cartoon. In the second experiment, recognizing the importance of diversity in generative models for design tasks, we use four models—FLUX.1-Schnell [50], Kandinsky 3.0 [51], PixArt-$\alpha$ [52], and Stable Diffusion XL—to generate images of the object “Sofa”. In a similar manner, we generate red cartoonish bird images using Kandinsky 3.0, Stable Diffusion 3 [53], and PixArt-$\alpha$. The increase in diversity is shown in Figure 1. The performance of our online algorithms and the comparison of samples generated by them with the baselines are shown in Figure 2. Finally, in the last experiment, we use Stable Diffusion XL to simulate models generating images of different dog breeds: Bulldog, German Shepherd, and Poodle, respectively. This illustrates the challenge of generating diverse object types with text-to-image models. Figure 5 illustrates the impact of using a mixture of models in the first and last experiments. The improvement in diversity is evident both visually and quantitatively, as reflected in the RKE Scores. As shown in Figure 6, our online algorithms consistently outperform others in generating more diverse samples.
@@
Figure 6: Performance comparison of online algorithms using the RKE metric for text-to-image generative models.
Text Generative Models
We utilize the OpenLLMText dataset [54], which comprises 60,000 human texts rephrased paragraph by paragraph using the GPT2-XL [55], LLaMA-7B [56], and PaLM [57] models. To extract textual features, we employ the RoBERTa Text Encoder. As shown in Figure 11(a) in Section C.3, the results demonstrate the advantage of our online algorithms, suggesting that our method applies not only to image generators, but also to text generators.
@@ 6.3 Optimal Mixture for Diversity and Quality via RKE and Precision/Density
Using RKE, we focus solely on the diversity of the models without accounting for their quality. To address this, we apply our methodology to both RKE and Precision [9], as well as RKE and Density [10]. We conduct experiments in which quality is a key consideration. We use four models: three are StyleGAN2-ADA models trained on the FFHQ dataset, each generating images with a truncation of 0.3 around randomly selected center points. The fourth model is StyleGAN2-ADA trained on CIFAR-10 [58]. The FFHQ dataset is used as the reference dataset, therefore, the model trained on CIFAR-10 acts as a low quality generator. Figures 7 and 12 demonstrate the ability of our algorithms to find optimal mixtures with a higher diversity/quality score.

### v2

@@ 6 Numerical Results
  • Mixture Oracle. In the mixture oracle algorithm (Section 5), an oracle tells us the optimal mixture $\bm{\alpha}^{*}$ in advance, and we pull arms randomly according to this distribution. The optimal mixture is calculated by solving the quadratic optimization in Section 4 on a large number of samples for each arm. The number of chosen samples varies based on the experiments. This is an unrealistic setting that only serves as a theoretical upper bound of the performance of any online algorithm. A realistic algorithm that performs close to the mixture oracle would be almost optimal.
@@
  • Successive Halving. The Success Halving algorithm (Karnin et al., 2013; Jamieson & Talwalkar, 2016; Chen & Ghosh, 2024) which serves as a second baseline for comparison.
@@
We used DINOv2-ViT-L/14 (Oquab et al., 2023) for image feature extraction, as recommended in (Stein et al., 2023), and RoBERTa (Liu et al., 2019) as the text encoder. Detailed explanation of the setup for each experiment is presented in Section 8.4.
@@ 6.1 Optimal Mixture for Diversity and Quality via KID
We conducted three experiments to evaluate our method using the Kernel Inception Distance (KID) metric. In the first experiment, we used five distinct generative models: LDM (Rombach et al., 2022), StyleGAN-XL (Sauer et al., 2022), Efficient-VDVAE (Hazami et al., 2022), InsGen (Yang et al., 2021), and StyleNAT (Walton et al., 2022), all trained on the FFHQ dataset (Karras et al., 2019). In the second experiment, we used generated images from four models [^9] FFHQ and LSUN-Bedroom datasets were downloaded from the dgm-eval repository (Stein et al., 2023) (licensed under MIT license): https://github.com/layer6ai-labs/dgm-eval.: StyleGAN (Karras et al., 2019), Projected GAN (Sauer et al., 2021), iDDPM (Nichol & Dhariwal, 2021), and Unleashing Transformers (Bond-Taylor et al., 2022), all trained on the LSUN-Bedroom dataset (Yu et al., 2015). This experiment followed a similar setup to the first. In the final experiment, we employed the truncation method (Marchesi, 2017; Karras et al., 2019) to generate diversity-controlled images centered on eight randomly selected points, using StyleGAN2-ADA (Karras et al., 2020), also trained on the FFHQ dataset. Figure 3 demonstrates that the mixture of generators achieves better KID scores compared to individual models. Additionally, the two Mixture-UCB algorithms consistently outperform the baselines.
@@ 6.2 Optimal Mixture for Diversity via RKE
We used the RKE Mode Count (Jalali et al., 2023) as an evaluation metric to show the effect of mixing the models on the diversity and the advantage of our algorithms Mixture-UCB-CAB and Mixture-UCB-OGD. The score in the plots is the RKE Mode Count, written as RKE for brevity.
@@
We conduct two experiments on diversity-limited generative models. First, we used eight center points with a truncation value of 0.3 to generate images using StyleGAN2-ADA, trained on the FFHQ dataset. In the second experiment, we applied the same model, trained on the AFHQ Cat dataset (Choi et al., 2020), with a truncation value of 0.4. As shown in Figure 7, the optimal mixture and our algorithms consistently achieve higher RKE scores. The increase in diversity is visually depicted in Figures 6 and 8.
@@
We used Stable Diffusion XL (Podell et al., 2024) with specific prompts to create three car image generators with distinct styles: realistic, surreal, and cartoon. In the second experiment, recognizing the importance of diversity in generative models for design tasks, we used five models—FLUX.1-Schnell (Labs, 2024), Kandinsky 3.0 (Vladimir et al., 2024), PixArt-$\alpha$ (Chen et al., 2023a), and Stable Diffusion XL—to generate images of the object “Sofa”. In a similar manner, we generated red bird images using Kandinsky 3.0, Stable Diffusion 3 (Esser et al., 2024), and PixArt-$\alpha$, as shown in Figure 2.. Finally, in the third experiment, we used Stable Diffusion XL to simulate models generating images of different dog breeds: Bulldog, German Shepherd, and Poodle, respectively. This illustrates the challenge of generating diverse object types with text-to-image models. Figure 10 demonstrates the impact of using a mixture of models in the first and third experiments. The improvement in diversity is evident visually and quantitatively, as shown by the RKE scores. Our online algorithms consistently generate more diverse samples than others, as illustrated in Figure 4.
@@
Figure 4: Performance comparison of online algorithms using RKE score of T2I generative models.
@@ 6.3 Optimal Mixture for Diversity and Quality via RKE and Precision/Density
Using RKE, we focus solely on the diversity of the arms without accounting for their quality. To address this, we apply our methodology to both RKE and Precision (Kynkäänniemi et al., 2019), as well as RKE and Density (Naeem et al., 2020). We conduct experiments in which quality is a key consideration. We use four arms: three are StyleGAN2-ADA models trained on the FFHQ dataset, each generating images with a truncation of 0.3 around randomly selected center points. The fourth model is StyleGAN2-ADA trained on CIFAR-10 (Krizhevsky et al., 2009). The FFHQ dataset is used as the reference dataset. Figures 5 and 13 demonstrate the ability of our algorithms in finding optimal mixtures with higher diversity/quality score.